# Ammann Bars for Octagonal Tilings

### v5

@@ 3.2 Finding fine projections
Of course with 4 variables and 16 equations there is no guarantee that a solution exists, and oftentimes when a projection respects the collinear condition in Definition 1 it is not valid: some tiles are superimposed in what should be a tiling. Figure 4 shows for instance what happens in the case of golden octagonal tilings (introduced in [4]) when the obtained matrix $A$ is used. To find a slope $E$ with a fine projection, we proceed as follows:
@@ 3.3 Defining the prototiles
Let $E$ be a $2$-dimensional irrational plane in $\mathbb{R}^{4}$ characterized by its subperiods and which admits a fine projection $\pi$. Consider a tiling with slope $E$ obtained using the fine projection $\pi$. Draw through each vertex of this tiling four lines directed by each of the projected subperiods $\pi(p_{i})$’s. Figure 8 shows what we obtain for a Cyrenaic tiling. These lines decorate the tiles of the tiling with segments that can take four different directions. All these decorated tiles, considered up to translation, define the wanted tileset. Note that the tileset does not depend on the initially considered tiling, because the $4\to 2$ tilings with a given irrational slope share the same finite patterns (this known fact is e.g. proven by Prop. 1 in [6]). We can now prove:
@@
Proof. We prove that the number of different intervals (distances) between two consecutive lines in a given direction is finite. This yields finitely many ways to decorate a tile by parallel segments, hence finitely many different tiles.
Consider a subperiod $p_{i}$ and the set $\mathcal{D}_{i}$ of all lines in $E$ directed by $\pi(p_{i})$ and passing through the vertices of the tiling, that is by all points $\pi(x)$ with $x\in\mathbb{Z}^{4}\cap(E+[0,1]^{4})$. Since the distance from a vertex to its neighbors is $||\pi(e_{k})||$ for some $k$, the interval between two consecutive lines of $\mathcal{D}_{i}$ is at most $d_{1}:=\max_{j\neq i}\{||\pi(e_{j})||\}$.
@@
Although the previous proof does not give an explicit bound on the number of tiles, it does give a constructive procedure to obtain these tiles. It is indeed sufficient to compute the constant $d$ (which depends on the subperiods and the projection), then to enumerate the $d$-maps (for example by enumerating all patterns of size $d$ and keeping only those which can be lifted in a tube $E+[0,1]^{4}$ – in practice we used a more efficient algorithm based on the notion of region [7] described in Appendix B – and, for each $d$-map, to draw the lines and enumerate the new decorated tiles obtained. In the case of Cyrenaic tilings, it is sufficient to enumerate the tiles which appear in the 5-atlas in terms of graph distance. We obtain 2 or 3 intervals in each direction, and the set $\mathcal{C}$ of 36 decorated prototiles in Figure 2.
@@ 4 Tiling with an FP-tileset
Proof. For more clarity, the diagram in Figure 10 summarizes the notations and the “traveling” between spaces that we use.
@@
In other words, no matter which copy of each tile yields the prototile in the tileset, for a given position $x$ in the prototile $\bar{x}$ will be the same. We also observe that $\bar{x}=0$ if and only if $x$ is a vertex of the tile. Indeed, if $\bar{x}=0$ then $\mu\in\mathbb{Z}$ because $\omega_{i}(p_{i})$ is an integer vector whose coordinates are relatively prime. And by definition $\bar{x}=\hat{v}+\mu\,\omega_{i}(p_{i})\bmod\mathbb{Z}^{3}$ where $v$ is a vertex of the tiling, so $\hat{v}$ is a vertex of the shadow and it follows that the $i$-shadow is $\omega_{i}(p_{i})$-periodic, thus $\bar{y}$ is also a vertex of the shadow, hence $y$ is a vertex of the tiling.
@@
The main result in [4] thus yields the following:
@@ B.1 Computing cut and project tilings
We use the duality between multigrids [10] and cut and project tilings. A grid is a set of regularly spaced parallel hyperplanes (lines in the case of an $n\to 2$ tiling). A multigrid is then a $d\times n$ matrix $G$ whose rows define the direction and spacing of each grid (normal vector to each hyperplane, with the norm giving the spacing), and a shift vector $S$ which specifies how each grid is translated away from the origin. The shift must be chosen such that no more than $d$ hyperplanes intersect in a point (this is generic).
@@ B.2 Finding the integer entries of a subperiod
Consider a slope $E$ generated by the vectors $u_{1},...,u_{d}\in\mathbb{R}^{n}$. We assume that the entries of the $u_{i}$’s are in $\mathbb{Q}(a)$ for some algebraic number $a$, because it has been proven in [7] to be a necessary condition to have local rules. We want to find a subperiod $p$ of $E$, that is a vector of $E$ with $d+1$ integer coordinates – the type of $p$ gives us the indices of its integer coordinates $a_{i_{1}},...,a_{i_{d+1}}$ ($1\leq i_{1}<i_{2}<...<i_{d+1}\leq n$). For this, we consider the $n\times(d+1)$ matrix $N$ whose columns are the $u_{j}$’s and the subperiod $p$ to be determined:
@@ B.6 Computing the r-atlas
Computing the $r$-atlas of a cut and project tiling mainly relies on the notions of window and region. We here briefly recall these notions; the interested reader can find more details in [7].

### v6

@@ 3.2 Finding fine projections
Of course with 4 variables and 16 equations there is no guarantee that a solution exists, and oftentimes when a projection respects the collinear condition in Definition 1 it is not valid: some tiles are superimposed in what should be a tiling. Figure 4 shows for instance what happens in the case of golden octagonal tilings (introduced in Bédaride and Fernique (2015a)) when the obtained matrix $A$ is used. To find a slope $E$ with a fine projection, we proceed as follows:
@@ 3.3 Defining the prototiles
Let $E$ be a $2$-dimensional irrational plane in $\mathbb{R}^{4}$ characterized by its subperiods and which admits a fine projection $\pi$. Consider a tiling with slope $E$ obtained using the fine projection $\pi$. Draw through each vertex of this tiling four lines directed by each of the projected subperiods $\pi(p_{i})$’s. Figure 8 shows what we obtain for a Cyrenaic tiling. These lines decorate the tiles of the tiling with segments that can take four different directions. All these decorated tiles, considered up to translation, define the wanted tileset. Note that the tileset does not depend on the initially considered tiling, because the $4\to 2$ tilings with a given irrational slope share the same finite patterns (this known fact is e.g. proven by Prop. 1 in Bédaride and Fernique (2017)). We can now prove:
@@
Proof.
We prove that the number of different intervals (distances) between two consecutive lines in a given direction is finite. This yields finitely many ways to decorate a tile by parallel segments, hence finitely many different tiles. Consider a subperiod $p_{i}$ and the set $\mathcal{D}_{i}$ of all lines in $E$ directed by $\pi(p_{i})$ and passing through the vertices of the tiling, that is by all points $\pi(x)$ with $x\in\mathbb{Z}^{4}\cap(E+[0,1]^{4})$. Since the distance from a vertex to its neighbors is $||\pi(e_{k})||$ for some $k$, the interval between two consecutive lines of $\mathcal{D}_{i}$ is at most $d_{1}:=\max_{j\neq i}\{||\pi(e_{j})||\}$.
@@
Although the previous proof does not give an explicit bound on the number of tiles, it does give a constructive procedure to obtain these tiles. It is indeed sufficient to compute the constant $d$ (which depends on the subperiods and the projection), then to enumerate the $d$-maps (for example by enumerating all patterns of size $d$ and keeping only those which can be lifted in a tube $E+[0,1]^{4}$ – in practice we used a more efficient algorithm based on the notion of region (Bédaride and Fernique, 2020) described in Appendix B – and, for each $d$-map, to draw the lines and enumerate the new decorated tiles obtained. In the case of Cyrenaic tilings, it is sufficient to enumerate the tiles which appear in the 5-atlas in terms of graph distance. We obtain 2 or 3 intervals in each direction, and the set $\mathcal{C}$ of 36 decorated prototiles in Figure 2.
@@ 4 Tiling with an FP-tileset
Proof.
For more clarity, the diagram in Figure 10 summarizes the notations and the “traveling” between spaces that we use.
@@
The main result in Bédaride and Fernique (2015a) thus yields the following:
@@ B.1 Computing cut and project tilings
We use the duality between multigrids de Bruijn (1981) and cut and project tilings. A grid is a set of regularly spaced parallel hyperplanes (lines in the case of an $n\to 2$ tiling). A multigrid is then a $d\times n$ matrix $G$ whose rows define the direction and spacing of each grid (normal vector to each hyperplane, with the norm giving the spacing), and a shift vector $S$ which specifies how each grid is translated away from the origin. The shift must be chosen such that no more than $d$ hyperplanes intersect in a point (this is generic).
@@ B.2 Finding the integer entries of a subperiod
Consider a slope $E$ generated by the vectors $u_{1},...,u_{d}\in\mathbb{R}^{n}$. We assume that the entries of the $u_{i}$’s are in $\mathbb{Q}(a)$ for some algebraic number $a$, because it has been proven in Bédaride and Fernique (2020) to be a necessary condition to have local rules. We want to find a subperiod $p$ of $E$, that is a vector of $E$ with $d+1$ integer coordinates – the type of $p$ gives us the indices of its integer coordinates $a_{i_{1}},...,a_{i_{d+1}}$ ($1\leq i_{1}<i_{2}<...<i_{d+1}\leq n$). For this, we consider the $n\times(d+1)$ matrix $N$ whose columns are the $u_{j}$’s and the subperiod $p$ to be determined:
@@ B.6 Computing the r-atlas
Computing the $r$-atlas of a cut and project tiling mainly relies on the notions of window and region. We here briefly recall these notions; the interested reader can find more details in Bédaride and Fernique (2020).